Find the value(s) of w so that the points (0,3) and (6,w) are 10 units apart.
Answer by Earlsdon(6103)

You can use the distance formula to solve this problem.
d = sqrt[(x2 - x1)^2 + (y2 - y1)^2], where: x1 = 0, x2 = 6, y1 = 3, y2 = w, and d = 10
10 = sqrt[(6-0)^2 + (w-3)^2]
10 = sqrt[36 + (w-3)^2] Square both sides.
100 = 36 + (w^2 - 6w + 9) Simplify.
w^2 - 6w + 45 = 100 Subtract 100 from both sides.
w^2 - 6w - 55 = 0 Solve the quadratic equation by factoring.
(w - 11)(w + 5) = 0 Apply the zero products principle.
w - 11 = 0, w = 11
w + 5 = 0, w = -5
The values of w are: 11 and -5